OpenStudy (arbyj):

how do i factor this polynomial; 9x^2+6x^3 helper will be rewarded !

OpenStudy (anonymous):

first what is common between the two?

OpenStudy (arbyj):

they're both divisible by 3

OpenStudy (anonymous):

3x^2 is common right?

OpenStudy (anonymous):

you group them and 3x^2(3+2x) I guess?

OpenStudy (arbyj):

i think so

OpenStudy (solomonzelman):

via rules of exponents: \(\color{#000000 }{ \displaystyle x^3=x^{2+1}=x^2\cdot x^1=x^2\cdot x }\) then, \(\color{#000000 }{ \displaystyle 9x^2+6x^3 }\) \(\color{#000000 }{ \displaystyle 3\cdot 3\cdot x^2+3\cdot 2 \cdot x^2 \cdot x }\) \(\color{#000000 }{ \displaystyle 3\cdot (3x^2)+ 2x \cdot (3x^2) }\) so the above ws correct

OpenStudy (solomonzelman):

the rule is: \(\color{#000000 }{ \displaystyle a\times b+a\times c=(b+c)a }\)

OpenStudy (solomonzelman):

if you think about it, it makes sense. Try calculating each side separately, \(\color{#000000 }{ \displaystyle 4\times 3+4\times 5=(3+5)4 }\)

OpenStudy (arbyj):

okay, thank you all !